Can someone provide the steps to Evaluate (16x)/1.4-x for x=0.35?
Answer by rothauserc(4718)

Evaluate (16x)/(1.4-x) for x=0.35
= (16 * 0.35) / (1.4 - 0.35)
= 5.6 / 1.05
= 5.33